The equation of a straight line passing through the point (-5,4) and which cuts an intercept of root(2) units between the lines x+y+1=0 and x+y-1 is?

Grade:12

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello Student. Please find the solutionx+y+1=0 and x+y-1 are two parallel lines and distance between them is mod|\frac{1-(-1)}{\sqrt{2}} |So distabce between them is 2/root2 = root2So our req line us perpendicular to these linesSo slope of required line is +1So eqn is (y-4)=1(x+5)thanks,